OpenStudy (anonymous):

Let f (x)=7x^(2)+2x+3. Find f(x+h)-f(x)

OpenStudy (anonymous):

First put in (x+h) in place of all the x's

OpenStudy (anonymous):

I have done this problem several times and keep getting it wrong

OpenStudy (anonymous):

[7(x+h)^2 + 2(x+h) +3]-(7x^2+2x+3)=

OpenStudy (anonymous):

I got 14hx+7h^2+2x+2h-2x

OpenStudy (anonymous):

and it was wrong

OpenStudy (anonymous):

7(x^2+2xh+h^2)+2x+2h-7x^2-2x-3=

OpenStudy (anonymous):

Like I said I keep doing it and I am not getting it right.. I think im burnt on this problem lol been over 30 minutes working on just this one

OpenStudy (anonymous):

7x^2+14xh+7h^2+2h-7x^2-3=

OpenStudy (anonymous):

14hx-3?

OpenStudy (anonymous):

7h^2+14xh-3

OpenStudy (anonymous):

if you miss it 3 times it gives you a new problem so i just missed it for the 3 rd time

OpenStudy (anonymous):

but thanks

OpenStudy (anonymous):

7h^2+14xh-2x-3. I left out a term